OpenStudy (anonymous):

Malory purchased a car for $21,825. It depreciates about 5.8% each year. What is the value of the car after six years?

OpenStudy (anonymous):

21,056.49

OpenStudy (anonymous):

21,065.49 * correction

OpenStudy (phi):

if you lose 5.8% value, you have 100-5.8 = 94.2 % value at the end of a year change that to a decimal: 94.2% is 0.942 now multiply your starting value 21,825 times 0.942 for each year that goes by

OpenStudy (nurali):

y = 21,825 - 0.058x x being years so... y = 21,825 - 0.058(6)